Twice the sum of four and a number is 36
2(4+x)=36
8+2x=36
2x=28
X=14

Found 2 solutions by stanbon, Ctrader:
Answer by stanbon(75887)

Twice the sum of four and a number is 36
------
2(4+x) = 36
4+x = 18
x = 14
------------
Cheers,
Stan H.

Answer by Ctrader(1)

Twice the sum of four and a number is 36
2(4+x)=36
8+2x=36
2x=28
X=14